For each of the following, prove that the line is a tangent to the parabola, and find the point of contact.
y=x-1/4 and y=x^2
I had simplified that to x^2-x+1/4
However, I cannot factorise that equation....in that form. I know that it MUST be a tangent; as when I used the discriminant, i.e. b^2-4ac; the answer was exactly 0.
Previously, with other questions I was able to clearly identify the value of x, by factorising; then, subsituting the value of x derived in this manner, into either of the original equations to find Y.
However, because I can't factorise here, I can't do that so am at a dead end.



Answer by rothauserc(4718)

x^2-x+1/4 = (x - (1/2))^2
x = 1/2 and y = 1/4 therefore
(1/2, 1/4) is the point of tangency